# GENIE: Watermarking Graph Neural Networks for Link Prediction

### v1

@@ Appendix B Statistical guarantee of non-trivial ownership in GENIE
The values of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ are given for $n=10$ different $\mathcal{M}_{clean}$ and ($\mathcal{D}_{wm}$, $\mathcal{M}_{wm}$) in TABLE B.1 (for SEAL), TABLE B.2 (for GCN), and TABLE B.3 (for GraphSAGE). The corresponding $p$-values are also mentioned. For each dataset and architecture, the $p$-value is observed to be below the significance level, i.e., $0.051-\tau=$ 0.05. Therefore, we reject $\mathbf{H}_{0}$ for each architecture and dataset as described in §4.3.
TABLE B.1: Non-trivial ownership results for SEAL.
@@
TABLE B.2: Non-trivial ownership results for GCN.
@@
TABLE B.3: Non-trivial ownership results for GraphSAGE.
@@
Appendix C Calculating watermark threshold
The watermark thresholds can be set either empirically or using statistics. We want to ensure a low FPR and FNR using statistical analysis. Therefore, the calculation of thresholds is not obvious given the non-normality of the sample data — $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$. There is also a dearth of the sample data as training multiple models is expensive. If there was no dearth, using the standard bootstrap hypothesis testing [66] would suffice to set the threshold. A direct application of standard bootstrap hypothesis testing would give thresholds marginally above, and in some cases even below the maximum of the different $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ s, due to the evident discreteness of the sample data. This estimate of the threshold can be used, but is prone to outliers specific to the sample data. We tackle this problem by drawing inspiration from the SBA [48].
The Bootstrap deals with data scarcity by resampling from the sample data multiple times and calculating the statistic based on this resampled data. Essentially, the Bootstrap is trying to capture the distribution of the statistic by resampling. SBA also does the same, but by resampling from a Kernel Density Estimate (KDE) [67] of the sample data instead of the sample data itself. This allows for values that are not in the sample data to be in the resampled data, making the Bootstrap smooth. We take inspiration from this and set our thresholds as follows:
  1. 1.
    Estimate the probability distribution of the sample data using KDE, with bandwidth set according to Silverman’s rule of thumb [53].
  2. 2.
    Approximate the population’s probability distribution by substituting it with the estimated probability distribution of the sample data.
  3. 3.
    Draw $n$ samples from the population’s probability distribution.
  4. 4.
    Find a threshold $t$ such that the number of false positives and false negatives are minimized.
If a threshold is found such that the FPR and FNR is 00 across multiple runs of our simulation, a statistical guarantee can be given that FPR and FNR are both less than $n^{-1}$. Figure C.1 plots the distribution of the data from TABLE II and calculates SEAL’s watermark threshold on the Yeast dataset. We use $n=10^{6}$ to ensure FPR and FNR to be less than $10^{-6}$.
Figure C.1: Population distribution of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ along with the corresponding watermark threshold $t$ with $n=10^{6}$. $\alpha$ and $\beta$ denote the FPR and FNR, while $\mu_{c}$ and $\mu_{w}$ denote the population mean of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$.
@@ Appendix D Additional results
Here, we present further results demonstrating the robustness of GENIE. Appendix D.1 lists the results when watermarking was applied to the GraphSAGE model while Appendix D.2 covers the results for the SEAL model. Here, we highlight the value of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ in red if it dropped below the watermark threshold (mention in TABLE III). Appendix D.3 presents additional results for non-ownership piracy test.
D.1 Robustness of GraphSAGE
TABLEs D.1-D.9 present results for different robustness tests for the GraphSAGE model.
@@ D.1 Robustness of GraphSAGE
D.2 Robustness of SEAL
TABLEs D.10-D.16 present results for different robustness tests for the SEAL model.
@@ D.2 Robustness of SEAL
D.3 Further results for non-ownership piracy
Due to the page length limit, we could only include results for one dataset for non-ownership piracy test in the main text. Now, we present the results for the other datasets in Figure D.1.
(a) C.ele
(b) USAir
(c) PPI
(d) Power
(e) arXiv
(f) Yeast

### v3

@@ Appendix C Statistical assurance of non-trivial ownership in Genie
The values of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ are given for $n=10$ different $\mathcal{M}_{clean}$ and ($\mathcal{D}_{wm}$, $\mathcal{M}_{wm}$) in TABLE C.1 (for SEAL), TABLE C.2 (for GCN), TABLE C.3 (for GraphSAGE), and TABLE C.4 (for NeoGNN). The corresponding $p$-values are also mentioned. For each dataset and architecture, the $p$-value is observed to be below the significance level, i.e., $0.051-\tau=$ 0.05. Therefore, we reject $\mathbf{H}_{0}$ for each architecture and dataset as described in §4.3.
TABLE C.1: Non-trivial ownership results for SEAL.
@@
TABLE C.2: Non-trivial ownership results for GCN.
@@
TABLE C.3: Non-trivial ownership results for GraphSAGE.
@@
TABLE C.4: Non-trivial ownership results for NeoGNN.
Dataset Model ($\mathcal{M}$) $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}}$ (%) $p$-value
Celegans $\mathcal{M}_{clean}$ 20.31 28.12 23.44 12.50 25.00 14.06 18.75 23.44 23.44 15.62 0.000
USAir $\mathcal{M}_{clean}$ 11.33 12.60 10.16 12.89 11.23 11.33 10.64 14.84 11.91 9.18 0.000
NS $\mathcal{M}_{clean}$ 23.11 20.22 28.89 26.44 31.78 26.44 17.33 28.89 23.11 31.78 0.000
Yeast $\mathcal{M}_{clean}$ 5.33 7.10 8.28 9.47 9.47 6.80 8.88 8.28 5.62 9.47 0.000
Power $\mathcal{M}_{clean}$ 50.00 45.00 45.00 50.00 50.00 50.00 50.00 50.00 50.00 45.00 0.000
$\mathcal{M}_{wm}$ 100.00 100.00 100.00 97.00 100.00 100.00 100.00 100.00 100.00 100.00
arXiv $\mathcal{M}_{clean}$ 8.38 1.17 9.42 7.64 1.25 8.15 9.01 8.20 7.05 9.86 0.000
$\mathcal{M}_{wm}$ 97.34 94.03 93.64 90.38 87.73 96.19 97.44 99.97 98.70 92.44
PPI $\mathcal{M}_{clean}$ 24.52 2.85 15.15 11.11 10.47 16.25 17.95 6.52 10.19 12.81 0.000
$\mathcal{M}_{wm}$ 100.00 99.91 97.43 97.61 95.32 94.63 97.52 97.06 100.00 96.97
@@ Appendix D Additional results
Here, we present further results demonstrating the robustness of Genie. Appendix D.1 presents additional results for non-ownership piracy test. Appendix D.2 lists results for differnt fine-pruning tests for the GCN model. Appendix D.3 lists the results when watermarking was applied to the GraphSAGE model while Appendix D.4 covers the results for the SEAL model. Here, we highlight the value of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ in red if it dropped below the watermark threshold (mention in TABLE III).
D.1 Non-ownership piracy
Genie injects watermark into an untrained model while $\mathcal{A}$ has access to only $\mathcal{M}_{wm}$. In a real-world setting, $\mathcal{A}$ can still generate her own pirated trigger set (using the method explained in § 4.1) and train stolen $\mathcal{M}_{wm}$ on the pirated trigger set to obtain $\mathcal{M}_{adv}$ (generally called a pirated model). Given that training on just the pirated trigger set might lead to decrease in $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $\mathcal{A}$ would want to identify an optimal number of epochs for training with the pirated trigger set such that $\mathcal{M}_{adv}$ has high AUC on both $\mathcal{D}_{test}$ and the pirated trigger set. Figure D.1 shows the variations in $\mathcal{M}_{adv}$’s performance on $\mathcal{D}_{wm}$, $\mathcal{D}_{test}$, and pirated trigger set during pirated watermark embedding process across different numbers of epochs for GCN over NS dataset; cf. Figure D.2 for other datasets.
Figure D.1: A representative example of $\mathcal{M}_{adv}$’s performance trajectory on $\mathcal{D}_{wm}$, $\mathcal{D}_{test}$, and pirated trigger set during embedding of pirated watermark across training epochs.
We see that $\mathcal{M}_{adv}$ performs well on $\mathcal{D}_{wm}$, $\mathcal{D}_{test}$, as well as on pirated trigger sets around $20^{th}$ epoch. If $\mathcal{O}$ challenges $\mathcal{A}$ to present her model at this point, $\mathcal{M}_{adv}$ will contain $\mathcal{A}$’s pirated watermark as well as $\mathcal{O}$’s watermark. However, $\mathcal{O}$ can present $\mathcal{M}_{wm}$ containing only her watermark. Thus, identifying the true owner will be easy in such a dispute. We further observe that around $250^{th}$ epoch, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ drops below the watermark threshold (cf. TABLE III), but $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$ falls below the tolerable utility loss (i.e., up to 10%; following the definition of failure in §5.1) $\mathcal{A}$ is willing to tolerate. Even if $\mathcal{A}$ chooses to train for even higher epochs while embedding the pirated watermark, $\mathcal{M}_{adv}$ continues to lose its utility; rendering the $\mathcal{M}_{adv}$ useless. Hence, we take the liberty to claim that Genie is robust against piracy attacks (i.e., $\mathcal{A}$ cannot fraudulently claim ownership or fabricate watermark over a pirated model). Now, we present the results for the other datasets in Figure D.2.
(a) C.ele
(b) USAir
(c) PPI
(d) Power
(e) arXiv
(f) Yeast
D.2 Fine-pruning GCN
TABLEs D.1-D.3 present results for different fine-pruning tests for the GCN model.
TABLE D.1: Impact of pruning + FTLL.
Dataset Prune Percentage (%)
$\text{No Pruning}^{\star}$ 20 40 60 80
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 93.75 90.62 81.25 79.68
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 91.21 90.42 87.50 69.82
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.77 99.77 98.88 97.55 93.11
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 91.71 90.53 85.20 92.30
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.00 99.00 95.00 87.00 72.00
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 86.51 81.81 59.24 30.07
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 96.32 88.52 79.43 51.97
Column with shows the values when $\mathcal{M}_{adv}$ = $\mathcal{M}_{wm}$.
TABLE D.2: Impact of pruning + RTLL.
Dataset Prune Percentage (%)
$\text{No Pruning}^{\star}$ 20 40 60 80
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 59.37 57.81 68.75 84.37
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 82.12 82.81 79.58 61.62
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.77 96.66 97.55 96.22 94.44
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 88.16 88.16 77.51 84.61
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.00 99.00 93.00 91.99 71.00
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 44.95 47.04 48.31 39.69
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 59.41 51.42 53.16 46.28
Column with shows the values when $\mathcal{M}_{adv}$ = $\mathcal{M}_{wm}$.
TABLE D.3: Impact of pruning + FTAL.
Dataset Prune Percentage (%)
$\text{No Pruning}^{\star}$ 20 40 60 80
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 71.87 60.93 78.12 62.50
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 81.49 73.19 76.31 63.28
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.77 89.11 84.66 88.66 92.22
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 100 100 99.40 81.65
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 99.00 88.00 87.00 84.00 67.99
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 54.45 48.18 33.05 17.00
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ (%) 100 78.60 75.39 66.20 45.36
Column with shows the values when $\mathcal{M}_{adv}$ = $\mathcal{M}_{wm}$.
D.3 Robustness of GraphSAGE
TABLEs D.4-D.12 present results for different robustness tests for the GraphSAGE model.
@@ D.3 Robustness of GraphSAGE
D.4 Robustness of SEAL
TABLEs D.13-D.19 present results for different robustness tests for the SEAL model.